# Coarse-grained entropy balance of the Szilard engine

### v2

@@ I Introduction
  • Many discussions of Maxwell’s demon work with information-theoretic terms. It has even been proposed to extend the second law to include terms relating to “information", see SU10 , SS15 . Concerning this question it has been objected EN99 , N25 that information-theoretic approaches are inadequate because a consistent theory for them is not yet available. We can leave this question open, because in our solution approach we only want to use physical terms in the narrower sense and not information-theoretical terms. We also include “Shannon entropy" and “mutual information" among the physical terms in the narrower sense, although this point could be debated. It could not be ruled out that our approach can also be interpreted in terms of information theory; but we do not need an extension of known physics to analyze Maxwell’s demon.
@@
  • As already mentioned, the paradoxical character of Szilard’s process can be seen in the fact that it represents a possible PM2, i. e. a cyclic process in which work is gained by reducing the temperature of a heat bath while leaving everything else unchanged. Such a process would reduce the total entropy of a closed system and thus violate the 2nd law. When it comes to the question of why the Szilard process does not represent a PM2 on closer inspection, the representatives of the various principles are largely in agreement: it is due to the altered state of the memory for the measurement result. The differences between the principles mentioned, however, arise from the entropy balance of the individual phases of the Szilard process, which we will therefore examine in more detail.
@@
On the other hand, every purely conditional action can be realized reversibly, i. e. by a permutation of the states of an extended system $\Sigma{\mathcal{A}}$. Although this is an almost trivial statement of set theory, the more challenging task is to realize the conditional effect through a physical model of an apparatus ${\mathcal{A}}$ and the interaction between $\Sigma$ and ${\mathcal{A}}$. For the Szilard engine, such a model is outlined in sections II and IV.
@@ II Szilard engine and discussion of various priciples
This (modified) Szilard process is therefore a candidate for a perpetuum mobile of the 2nd kind (PM2). This term usually refers to a thermodynamic cyclic process in which heat is completely converted into work, whereby the overall state of the system + environment does not change, except for a small temperature reduction of the heat bath from which the converted heat originates and the content of an external energy store that absorbs the work gained. During a cycle of a PM2, the entropy of the system would therefore remain constant, while the entropy of the heat bath would decrease.
Thus, a PM2 violates the second law. The impossibility of a PM2 is often regarded as an equivalent formulation of the 2nd law that does not require an explicit entropy definition. We can leave the question of equivalence here, since we are only interested in the question of whether the Szilard process violates the 2nd law or not. In this context, arguments concerning PM2 and arguments concerning the entropy balance are usually mixed, also in S29 . We will try to separate these two types of arguments.
@@
We only consider the system consisting of the container, the one-molecule gas and the memory; the rest of the world is referred to as “environment". Then we claim the following results on the entropy changes $\Delta H$ of the various phases of the Szilard process, see Figure 1. $\Delta H$ is always understood as $\Delta H=H_{\scriptstyle{fin}}(\ldots)-H_{\scriptstyle{ini}}(\ldots)$ (with self-explaining notation) and we refer to it as an increase if $\Delta H>0$ and a decrease if $\Delta H<0$. Entropy is considered dimensionless by setting $k_{\rm B}=1$, although the Boltzmann constant will appear explicitly in some formulas in the usual way. For the following we assume that the concept of mutual information is known or refer to its later definition in subsection III.1.
@@ III.2 Conditional actions
In the following subsections, we will present the theory of pure condition actions in some detail, as these are easier to understand. In particular, a pure conditional action will always decrease the entropy of the system or keep it constant. A general transformation of the kind (9) can increase or decrease the entropy, and hence the non-increase of entropy has to be additionally postulated for non-pure conditional actions. However, non-pure conditional actions play a crucial role in our realization of the Szilard engine, which is why we also have to deal with the general case.
@@ III.3 Pure conditional actions
Let $f:\Sigma\to\Sigma$ be a pure conditional action, $p_{i}>0$ for all $i\in\Sigma$, and (11) be the corresponding transformation of probabilities. Then $H(p^{\prime})\leq H(p)$ and $H(p^{\prime})=H(p)$ iff $f$ is a bijection.
The proof can be found in S20 , section 5. For the convenience of the reader we will reproduce it here with slight variations. If $f$ is a bijection then the Shannon entropy is unchanged, $H(p^{\prime})=H(p)$. If $f$ is not a bijection then it can be written as a finite composition of maps which map exactly two states onto the same one and are injective else. Thus it suffices to prove $H(p^{\prime})<H(p)$ for maps $f$ satisfying, say, $f(1)=f(2)=j$ and $f$ being injective on the set $\{3,4,\ldots,n\}$ (the definition domain of $f$ consists of $n$ states with positive probability which may be less than $N=|\Sigma|$). Note that $p^{\prime}(j)=p(1)+p(2)$ and $p^{\prime}(f(i))=p(i)$ for $i=3,4,\ldots,n$. It follows that
@@
which holds due to $p(i)>0$ for all $i=1,\ldots,n$ and $\log$ being a strictly increasing function. $\Box$
@@ III.6 Models for general conditional actions and modified S principle
Additionally we will postulate that $H(p^{\prime})\leq H(p)$ since this is not automatically satisfied in thee general case.
@@ IV.1 General description
When the pendulum has reached its lowest point, it has an approximate kinetic energy of $E_{\scriptstyle{kin}}=q\,E\,a=\frac{1}{2}\theta_{y}\omega^{2}$ and a $y$-component of the angular momentum $L_{y}=\theta_{y}\omega$, where $\omega$ denotes its angular velocity. We place another auxiliary particle with mass $M=\theta_{y}/{a^{2}}$ in the plane $z=-a$ so that the pendulum collides with it elastically at its lowest point and transfers its kinetic energy and angular momentum to the auxiliary particle. After the collision, the pendulum comes to a standstill (except for small thermally excited oscillations) and the auxiliary particle is assumed to move frictionless along the line $z=-a,y=-a/2$ with the speed $v$, which, after a short calculation, is given by
$v=a^{3/2}\sqrt{\frac{2qE}{\theta_{y}}}\;.$ (38)
If the molecule initially was in the lower box, the auxiliary particle would remain at rest. Thus the state of the auxiliary particle after the conditional action can be viewed as a record of the position measurement of the molecule. Not that in contrast to the general description in Section II the memory has only two states since the state 00 (the auxiliary particle being at rest) is also used as the default value. But this does not change the entropy balance.
@@ IV.2 Entropy balance
We next calculate the entropies. The symbol $\Sigma$ refers to the one-particle gas, and ${\mathcal{A}}$ to the remaining parts of the Szilard engine, consisting of the auxiliary particle ${\mathcal{A}}_{\scriptstyle{aux}}$ and the remainder ${\mathcal{A}}_{\scriptstyle{r}}$. We obtain
@@ IV.3 Entropy balance of a modified Szilard engine
At the beginning of the process, two partitions are inserted this time, which divide the container into three equal partial volumes, say, into three cubes with edge length $a$. As for the two-boxed Szilard engine we will apply an electric field that accelerates the charged molecule downwards. Only when the molecule is in the upper box will the Szilard pendulum tilt to the left and, as in the previous case, hit the auxiliary particle. The state of the auxiliary particle after the conditional action thus represents the result of the measurement of whether the molecule was originally either in the upper box or in one of the other two boxes; this is a measurement with two possible outcomes, not three. After the pendulum has come to a standstill, we obtain a state in which the molecule is in the lower box with a probability of $2/3$ and in the middle box with a probability of $1/3$. In contrast to the two-boxed engine, this state is still correlated with the state of the auxiliary particle with $p(v_{0})=2/3$ and $p(v_{1})=1/3$.
@@
In fact, a short calculation shows that by pulling out the lower partition wall, the entropy of the gas + auxiliary particle system $\Sigma\,{\mathcal{A}}_{\scriptstyle{r}}$ increases by the amount $\Delta\widetilde{H}=\frac{1}{3}\log 2$, taking into account the vanishing mutual information at the end of the process. On the other hand, due to the isothermal expansion the entropy of the heat bath decreases by the amount $\Delta H^{\prime}=\log 3/2$, which would have to be interpreted as the entropy loss of a cycle of the alleged PM2. Obviously,

### v3

@@ I Introduction
  • Many discussions of Maxwell’s demon work with information-theoretic terms, see, e. g., MNV09 . It has even been proposed to extend the second law to include terms relating to “information", see SU10 , SS15 . Concerning this question it has been objected EN99 , N25 that information-theoretic approaches are inadequate because a consistent theory for them is not yet available. We can leave this question open, because in our solution approach we only want to use physical terms in the narrower sense and not information-theoretical terms. We also include “Shannon entropy" and “mutual information" among the physical terms in the narrower sense, although this point could be debated. It could not be ruled out that our approach can also be interpreted in terms of information theory; but we do not need an extension of known physics to analyze Maxwell’s demon.
@@
  • As already mentioned, the paradoxical character of Szilard’s process can be seen in the fact that it represents a possible PM2, i. e. a cyclic process in which work is gained by extracting heat from a heat bath while leaving everything else unchanged. Such a process would reduce the total entropy of a closed system and thus violate the 2nd law. When it comes to the question of why the Szilard process does not represent a PM2 on closer inspection, the representatives of the various principles are largely in agreement: it is due to the altered state of the memory for the measurement result. The differences between the principles mentioned, however, arise from the entropy balance of the individual phases of the Szilard process, which we will therefore examine in more detail.
@@
On the other hand, every pure conditional action can be realized reversibly, i. e. by a permutation of the states of an extended system $\Sigma{\mathcal{A}}$. Although this is an almost trivial statement of set theory, the more challenging task is to realize the conditional effect through a physical model of an apparatus ${\mathcal{A}}$ and the interaction between $\Sigma$ and ${\mathcal{A}}$. For the Szilard engine, such a model is outlined in sections II and IV.
@@ II Szilard engine and discussion of various priciples
This (modified) Szilard process is therefore a candidate for a perpetuum mobile of the 2nd kind (PM2). This term usually refers to a thermodynamic cyclic process in which heat is completely converted into work, whereby the overall state of the system + environment does not change, except for negligible decrease in temperature of the heat bath from which the converted heat originates and the content of an external energy store that absorbs the work gained. During a cycle of a PM2, the entropy of the system would therefore remain constant, while the entropy of the heat bath would decrease.
Thus, a PM2 violates the second law. The impossibility of a PM2 is often regarded as an equivalent formulation of the 2nd law that does not require an explicit entropy definition.
We can leave the question of equivalence open here, because we are only interested in whether the Szilard process violates the 2nd law or not. In this context, arguments concerning PM2 and arguments concerning the entropy balance are usually mixed, also in S29 . We will try to separate these two types of arguments.
@@
For the entropy balance, it is important to define the system boundaries precisely. For our approach, the following definition is useful: The system (later referred to as $\Sigma{\mathcal{A}}$) comprises the one-molecule gas, the memory and the entire setup required to operate the Szilard process, including the heat bath necessary for isothermal expansion. The rest of the world is referred to as “environment"; it contains the energy reservoir to absorb the work done in the Szilard process. In addition, entropy is released to the environment when the memory is erased, i. e. possibly to a second heat bath, see section IV.1.
Then we claim the following results on the entropy changes $\Delta H$ of the various phases of the Szilard process, see Figure 1. $\Delta H$ is always understood as $\Delta H=H_{\scriptstyle{fin}}(\ldots)-H_{\scriptstyle{ini}}(\ldots)$ (with self-explaining notation) and we refer to it as an increase if $\Delta H>0$ and a decrease if $\Delta H<0$. Entropy is considered dimensionless by setting $k_{\rm B}=1$, although the Boltzmann constant will appear explicitly in some formulas in the usual way. For the following we assume that the concept of mutual information is known or refer to its later definition in subsection III.1.
@@ III.2 Conditional actions
In the following subsections, we will present the theory of pure conditional actions in some detail, as these are easier to understand. In particular, a pure conditional action will always decrease the entropy of the system or keep it constant. A general transformation of the kind (9) can increase or decrease the entropy, and hence the non-increase of entropy has to be additionally included in the definition of non-pure conditional actions. However, non-pure conditional actions play a crucial role in our realization of the Szilard engine, which is why we also have to deal with the general case.
@@ III.3 Pure conditional actions
Let $f:\Sigma\to\Sigma$ be a pure conditional action and (11) be the corresponding transformation of probabilities. Then $H(p^{\prime})\leq H(p)$ and $H(p^{\prime})=H(p)$ iff $f$ is a permutation if restricted to $\Sigma_{+}^{(p)}:=\{i\in\Sigma\left|\right.p(i)>0\}$.
The proof can be found in S20 , section 5. For the convenience of the reader we will reproduce it here with slight variations.
Let us first assume $\Sigma=\Sigma_{+}^{(p)}$. If $f$ is a bijection then the Shannon entropy is unchanged, $H(p^{\prime})=H(p)$. If $f$ is not a bijection then it can be written as a finite composition of maps which map exactly two states onto the same one and are injective else. Thus it suffices to prove $H(p^{\prime})<H(p)$ for maps $f$ satisfying, say, $f(1)=f(2)=j$ and $f$ being injective on the set $\{3,4,\ldots,n\}$ (the definition domain of $f$ consists of $n$ states with positive probability which may be less than $N=|\Sigma|$). Note that $p^{\prime}(j)=p(1)+p(2)$ and $p^{\prime}(f(i))=p(i)$ for $i=3,4,\ldots,n$. It follows that
@@
which holds due to $p(i)>0$ for all $i=1,\ldots,n$ and $\log$ being a strictly increasing function.
Returning to the general case we first note that $H(p^{\prime})\leq H(p)$ remains valid if the limit process $p_{i}\to 0$ is carried out for $i\in\Sigma_{0}^{(p)}:=\{i\in\Sigma\left|\right.p(i)=0\}$. However, from $H(p^{\prime})=H(p)$ we can only conclude that $f$ acts as a permutation on $\Sigma_{+}^{(p)}$ because the action of $f$ on $\Sigma_{0}^{(p)}$ does not change the value of $H(p^{\prime})$. $\Box$
@@ III.6 Models for general conditional actions and modified S principle
Let $p$ denote the mixed state with maximal entropy, i. e., $p(i)=1/N$ for all $i\in\Sigma$. Then the stochastic matrix $T$ is doubly stochastic iff $Tp=p$. Hence, if $T$ is not doubly stochastic, then $Tp\neq p$ and hence $H(Tp)<H(p)$ for this special state $p$. But generally, it cannot be excluded that $T$ increases the entropy for certain states $p$. As a physical example we note that the contact of a system with a heat bath with temperature $\theta$ (without assuming complete thermalization) can be described by a stochastic matrix $T$, see SSG21 . Then we would obtain $H(Tp)<H(p)$ if the temperature of the system exceeds $\theta$ and the system is cooled down by the contact, while $H(Tp)>H(p)$ in the other case if the system’s temperature is less then $\theta$.
Therefore, in the definition of conditional actions, we will additionally require that $H(p^{\prime})\leq H(p)$ holds, as this is not automatically fulfilled in the general case of stochastic matrices.
@@ IV.1 General description
This condition also implies that in thermal equilibrium the spatial probability density will be concentrated at the bottom of the box in which the molecule is located. This is in sharp contrast to the case of pure gravity, where the spatial probability density within the box is practically constant.
When the pendulum has reached its lowest point, it has an approximate kinetic energy of $E_{\scriptstyle{kin}}=q\,E\,a=\frac{1}{2}\theta_{y}\omega^{2}$ and a $y$-component of the angular momentum $L_{y}=\theta_{y}\omega$, where $\omega$ denotes its angular velocity.
We place another auxiliary particle with mass $M$ in the plane $z=-a$ so that the pendulum collides with it elastically at its lowest point and after the collision the auxiliary particle moves frictionless along the line $z=-a,y=-a/2$ with the speed $v$. In order that the pendulum completely transfers its kinetic energy and angular momentum to the auxiliary particle the following equations have to be fulfilled:
$q\,E\,a=\frac{1}{2}\theta_{y}\omega^{2}=\frac{M}{2}v^{2},\quad\theta_{y}\omega=M\,v\,a\;.$ (38)
The positive solutions to these equations read
$M=\frac{\theta_{y}}{a^{2}},\quad v=a^{3/2}\sqrt{\frac{2qE}{\theta_{y}}},\quad\mbox{and }\omega=\sqrt{\frac{2qaE}{\theta_{y}}}\;.$ (39)
We will choose the mass of the auxiliary particle correspondingly.
If the molecule initially was in the lower box, the auxiliary particle would remain at rest. Thus the state of the auxiliary particle after the conditional action can be viewed as a record of the position measurement of the molecule. Note that in contrast to the general description in Section II the memory has only two states since the state 00 (the auxiliary particle being at rest) is also used as the default value. But this does not change the entropy balance.
@@
In our approach, we have to consider the heat bath as a finite system, while the usual idealization of heat baths makes the transition to infinite systems. This is not a real problem, as we will also make analogous approximation assumptions concerning the state of the heat bath before and after the Szilard process.
@@ IV.2 Entropy balance
We next calculate the entropies.
The symbol $\Sigma$ refers to the one-molecule gas, and ${\mathcal{A}}$ to the remaining parts of the Szilard engine including the heat bath, as explained in Section II. Part of ${\mathcal{A}}$ consists of the auxiliary particle denoted by ${\mathcal{A}}_{\scriptstyle{aux}}$, which acts as a memory, while all the rest of ${\mathcal{A}}$ is captured by the symbol ${\mathcal{A}}_{\scriptstyle{r}}$.
We obtain
@@ IV.3 Entropy balance of a modified Szilard engine
At the beginning of the process, two partitions are inserted this time, which divide the container into three equal partial volumes, say, into three cubes with edge length $a$.
As in the Szilard machine with two boxes, an electric field is applied which exerts a downward force on the charged molecule and can cause the Szilard pendulum to rotate macroscopically. But this only happens if the molecule is in the upper box. Only then will the pendulum tilt to the left and, as in the previous case, hit the auxiliary particle. If the molecule is in the middle box the pendulum will remain in a stable rest position because, due to the strong electric field, the molecule will be found near the bottom of the box with overwhelming probability, see the analogous remark in subsection IV.1.
The state of the auxiliary particle after the conditional action thus represents the result of the measurement of whether the molecule was originally either in the upper box or in one of the other two boxes; this is a measurement with two possible outcomes, not three. After the pendulum has come to a standstill, we obtain a state in which the molecule is in the lower box with a probability of $2/3$ and in the middle box with a probability of $1/3$. In contrast to the two-boxed engine, this state is still correlated with the state of the auxiliary particle with $p(v_{0})=2/3$ and $p(v_{1})=1/3$.
@@
In fact, a short calculation shows that by pulling out the lower partition wall, the entropy of the gas + auxiliary particle system $\Sigma\,{\mathcal{A}}_{\scriptstyle{aux}}$ increases by the amount $\Delta\widetilde{H}=\frac{1}{3}\log 2$, taking into account the vanishing mutual information at the end of the process. On the other hand, due to the isothermal expansion the entropy of the heat bath decreases by the amount $\Delta H^{\prime}=\log 3/2$, which would have to be interpreted as the entropy loss of a cycle of the alleged PM2. Obviously,